OpenStudy (anonymous):

verify the identity cos^2 2x=1-4cos^2x sin^2x any help would be greatly appreciated ive been on this question for over 30 minutes i've tried using cos2theta double angle formulas but im having difficulty

OpenStudy (loser66):

from the right handside \[1-4cos^2(x)sin^2(x) = 1-(2cos(x)sin(x))^2\\=1-sin(2x)^2 = cos^2(2x)\] Dat sit

OpenStudy (loser66):

Yes or No?

OpenStudy (anonymous):

Thank you so much!